OpenStudy (amberlykhan):

A sports sold 200 tennis racquets. Some were sold at $18 each and the rest were sold at $33 each. The total receipts fro these sales were $4,800. How many racquets did the store sell at $18 each?

OpenStudy (samanthagreer):

Alright you know that there are two types of raquets (lets name them x and y shall we,) you also know that the total sold is 200 raquets and $4800 was made. Set up the system: 18x + 33y = 4800 ($18 and $33 being the prices, and 4800 the total) x + y = 200 (total sold) Now just solve for x, as it represents the $18 raquets. 33x+18(200-x)=4800 33x+3600-18x=4800 15x=1200 so now go ahead and tell me what 1200 divided by 15 is

OpenStudy (amberlykhan):

im soooo confused... howd u solve for x, howd u get those equations

OpenStudy (amberlykhan):

@samanthagreer

OpenStudy (samanthagreer):

ok so there were 200 sold coorect and there were $4800 made correct

OpenStudy (amberlykhan):

yes

OpenStudy (samanthagreer):

ok so lets set it up we have 18x + 33y = 4800 meaning that some were sold for 18 and the x means how many were sold but we dont know so that is what we are finding and the same thing with 33 and y and what do you do when a varibale or a number is right next to another number with out a sign

OpenStudy (samanthagreer):

do you multiply it

OpenStudy (amberlykhan):

yes

OpenStudy (amberlykhan):

but i dont see that

OpenStudy (samanthagreer):

great job ye you would multipy it so after that we have this ($18 and $33 being the prices, and 4800 the total) x + y = 200 (total sold) Now just solve for x, as it represents the $18 raquets.

OpenStudy (samanthagreer):

are you understanding

OpenStudy (amberlykhan):

im just confused how would u solve for x since there are two seperate equations and two variables in both equations

OpenStudy (samanthagreer):

ok well we are getting there but from what i have already told you do you understadn that part

OpenStudy (amberlykhan):

yes

OpenStudy (samanthagreer):

great so now we will putting this together ok so i have these equations 33x+18(200-x)=4800 33x+3600-18x=4800 15x=1200 now tell me from this how do we find x 15x=1200

OpenStudy (samanthagreer):

if you dont know just tell me and ill explain

OpenStudy (amberlykhan):

i just dont get how u got those equations

OpenStudy (samanthagreer):

well im not sure how to explain that but how about we finish going through the problem and then you might understand

OpenStudy (amberlykhan):

ok

OpenStudy (amberlykhan):

my firend showed me this but idk

OpenStudy (amberlykhan):

i think thats how im supposed to solve it

OpenStudy (samanthagreer):

ok so now tell me what is 1200 divided by 15

OpenStudy (amberlykhan):

can u explain that one

OpenStudy (amberlykhan):

80

OpenStudy (amberlykhan):

@samanthagreer

OpenStudy (samanthagreer):

ok so since we have that then that means that x=80 sold at $33 so we know that part but we still need to find the 18 dollar ones not tell me agina what x =

OpenStudy (amberlykhan):

80

OpenStudy (samanthagreer):

ok so solve this 200-x=

OpenStudy (amberlykhan):

120

OpenStudy (amberlykhan):

120 raquets sold at $18 each

OpenStudy (amberlykhan):

but im soooooo sorry im still confused

OpenStudy (samanthagreer):

200-x=120 sold at $18 well that is the answer i wil put everything agian in case you want to look over is agianAlright you know that there are two types of raquets (lets name them x and y shall we,) you also know that the total sold is 200 raquets and $4800 was made. Set up the system: 18x + 33y = 4800 ($18 and $33 being the prices, and 4800 the total) x + y = 200 (total sold) Now just solve for x, as it represents the $18 raquets. 33x+18(200-x)=4800 33x+3600-18x=4800 15x=1200 x=80 sold at $33 200-x=120 sold at $18

OpenStudy (amberlykhan):

can u explain what my friend did

OpenStudy (samanthagreer):

im not sure that was how i have always been taght to do it sorry

OpenStudy (amberlykhan):

its okay thanks for taking ur time to help me i really appreciate it and i will look over the work too see if i get it

OpenStudy (samanthagreer):

ok i hope you are able to understand

OpenStudy (samanthagreer):

please remember to close your question